using matlab, write a script for the following that prompts the user to input the amount of cases bought as a vector, calculates the cost, and stores the correspondind cost in a row beneath the demand:

the price of produce is calculated as follows:

if 30 cases of oranges are bought, the cost is 4 dollars per case

if more than 30, but less than 100 cases are bought, the cost is 10 for the first 30 cases, then 3 dollars for every case in excess of 30.

if more than 100 cases are bought, the cost is 5 dollars for the first 100 cases bought plus 1 dollar for every case in excess of 100.

a basic fee of 2 dollars is charged regardless of how many cases are bought.
Answer #1
cases = zeros(2, 1); % 2x1 marix first row contains demand and second row contains cost
cases(1) = input('Enter amount of cases bought');

cost = 0;
if cases(1) > 100
    cost = 5*100; % 5 dollar per case for first 100 cases
    cost = cost + 1*(cases(1) - 100); % 1 dollar for every case in excess 100
elseif cases(1) > 30
    cost = 10*30; % 10 dollar per case for first 30 cases
    cost = cost + 3*(cases(1) - 30); % 3 dollar for every case in excess 30
elseif cases(1) == 30
    cost = 4*cases(1); % the cost is 4 dollars per case
end
cost = cost + 2; % 2 dollars charged regardless of number of cases bought

cases(2) = cost; % Store in row beneath the demand

SAMPLE OUTPUT:

COMMAND WINDOW >>final Enter amount of cases bought >> cases Cases 377

FOR ANY HELP JUST DROP A COMMENT